18)
solve the equation: tan^2 3x=3

Answer by stanbon(75887)

solve the equation: tan^2 (3x) = 3
tan(3x) = sqrt(3) or tan(3x) = -sqrt(3)
3x = pi/3 or 3x = (4/3)pi ; 3x = (2/3)pi or (5/3)pi
x = pi/9 or x = (4/9)(pi ; x = (2/9)pi or (5/9)pi
===========================
Cheers,
Stan H.